BCAL Reconstruction Meeting 2015-03-19
Video Conferencing Information
Meeting Time: 11:00 a.m. EDT
- To join via a Web Browser, go to the page  https://bluejeans.com/907185247.
- To join via Polycom room system go to the IP Address: 22.214.171.124 (bjn.vc) and enter the meeting ID: 907185247.
- To join via phone, use one of the following numbers and the Conference ID: 907185247.
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
- More information on connecting to bluejeans is available.
Participant Direct Lines
- JLab Phone: in CC F326 is 757-269-6460 (usual room)
- JLab Phone in CC L207 is 757-269-7084
- Phone in the Regina Video-conference Suite is 306-585-4204
- Check new dark hits code after David's adjustments (Tegan)
- Shower curvature tables: compare KLOE to IU codes: check gap in cluster handling between codes (Tegan, Andrei)
- z-coordinate determination from up/down amplitude ratio (Tegan, Andrei)
- Cleanup: (Tegan, Will L.)
- Set variable “failures” to a number other than zero (e.g. -9999 or similar)
- Comb code for constants to be pulled out into a database
- Sampling fraction tables
- Action Items
- Commissioning Status Quo
- Shower definitions, merging and inclusion of single-ended hits (Mark, Elton)
- pi0 BCAL threshold and gain factors (Will M.)
- Shower Curvature (Tegan, Andrei)
- BDT effort (Tegan/Justin)
- Any other business
Attendees: Mark D., Elton, Will M. (JLab), Will L. (CMU), Noemi, Tegan, Zisis (Regina)
- Elton announced the first CW beam. Check on rad levels. Accelerator will be opened for 4-6 hours. Plan is to run through the weekend. ## The solenoid had a hiccup (trip) at 600 A. The quench detector fired without there actually being a quench. There was a spike in the voltage of the subcoils and pickup coils. The magnetic field did change. The cause of this hiccup is unknown.
- Action Items: Tegan is looking into the dark hist code.
- Commissioning Status Quo
- Shower definitions, merging and inclusion of single-ended hits (Mark, Elton) See Log 3327412 Elton talked about the clusters seen in the examples of this log entry. Two green circles are seen instead of the expected one. All hits are assumed to be photons/shower particles, however, here Elton used charged particles which curve (sideways) coming into the BCAL. Parameters were changed in the cluster algorithm; clusters are together in (phi, theta, time). The windows around the parameters are nominally 5 sigma. A surprise in this work was that not all hist are included in a shower and single ended hits are also not included. Mark mentioned that there is no code for this at present but will be in the future. Will L. said that it is not clear that clusters should be merged, after all the code is optimized for photons and not charged particles. Elton said that both showers are matched to a track, but the code is Will M.'s and does not have the standard matching. The efficiency in the reconstruction varies from 70-80% up to 95%, the latter if single-ended hits are included. Note that the BCAL positional information is not used yet, and as a result the tracks end up in the wrong place. The BCAL display button shows the up and downstream hits separately. Single ended hits should result in gaps in the Points but that is not seen here.
- pi0 BCAL threshold and gain factors (Will M.) Will looked at the pi0 signal over background versus E in the events, to decide where our spring running trigger should be. Nothing revealing out of this so far. It looks like the fall running was 1.5 GeV. Mark wondered why it is so high. The range for the gain factors was 0.02-0.08. Why so large? Investigation is still ongoing. Elton mentioned that the light guide acceptances affect gain factors, and the latter also depend on the attenuation length indirectly (factor x2 variation in this), but the gain factor range is x4. Mystery. Will M. will update the constants in the ccdb today. There was some discussion on the possible reflections at the end of the calorimeter.
- Tegan is working on the 'hot' scaler channels. Those have been there for a long time. The thresholds set are ~2.5MV in the ADCs and 30 mV (which corresponds to ~6mV in the ADCs) in the discriminators. The latter could be taken to 35 mV to see if there is a difference in the 'hot' channels. Maybe those channels are slightly noisier or their baseline is different.
- Shower Curvature (Tegan, Andrei)Tegan is looking into a new way of 'making' showers. Right now, the IU code uses (E, theta) and Andrei's translation tables are then used to get the z-bins of the shower parts. Then the code loops back and attempts to put the hits into the right showers. One issue is that the Points in shower reconstructs outside the BCAL (~420 cm) which then forces the cluster to be inside the BCAL, so that Andrei's code does not see the them. 50% of the Points are seen. Will L. commented that the code handles the outside clusters by 'pushing' them towards the centre of the BCAL, which does introduce a bias.
- BDT effort (Tegan/Justin): progress
- Any other business: none.